August 26, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, August 26, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .


  1. Announcements
  2. Summary of Detector Performance
  3. Simulations
    • Generated Files
      • Detcom 02 - /cache/mss/halld/detcom/02/rest
      • DC2 - /mss/halld/data_challenge/02
    • New Simulation Plans
  4. Subdetector Reports
  5. Calibration Processing
  6. AOB


Attending: Sean (NU); Simon, Adesh, Paul, Kei, Alex B., Luke, Justin, Elton, Mark I. (JLab); Curtis, Will M., Naomi, Mike S. (CMU); Eric, Mahmoud (FIU); Brad (FSU)

  1. Announcements
    • The next offline monitoring run will be this weekend, so please get your requests in by noon on Friday. The plan is to run over the Spring and Summer running periods, to also analyze the recent cosmic runs that were taken. This run will be processed using Kei's "classic" scripts, since SWIF is currently down and its maintainer is on vacation.
  2. Summary of Detector Performance
    • Curtis has received updates from most detector groups, FCAL being a notable exception. He will contact the beamline and online groups to fill in their sections.
  3. Simulations
    • The Detcom 02 round of simulations has finished, and are accessible at the directory mentioned in the agenda. More details on the simulations, including how to analyze them, can be found on the linked wiki pages. Standard monitoring plugins were also run over these simulations, and the results can be found on the offline monitoring pages.
    • There was a discussion about further simulation plans
      • The next planned simulations are for one run of the spring simulations (1300 A solenoid, coherent bremsstrahlung) to be generated with the updated BCAL simulation code. Mark I. will tag a new sim-recon version first.
      • It would also be useful to generate some data corresponding to projected fall running conditions. Proposal: 12 GeV e- beam, 1300 A solenoid magnetic field, coherent bremsstrahlung from 20um diamond radiator. Variations with the coherent peak on and off the microscope are possible.
      • A large scale DC4 is possible in the Jan.-Feb. 2016 range. We might want to wait for HDGeant4 first, though.
  4. Subdetector Reports
    • FCAL - Adesh has studied the effect of using event-by-event pedestals and average pedestals stored in the CCDB. The average pedestals give a better pi0 mass and smaller width of the peak, so they are preferred. He continues to study simulations.
    • BCAL - Work progresses on the BCAL TDCs. Mike S. has added some timing and timewalk constants based on Run 2931. The BCAL reconstruction code uses ADC times by default, so this work can continue without affecting analysis.
    • CDC - Simon and Mike S. are still working on the "sagging straw" problem. Mike is looking at cosmic data taken recently - there is about 2 weekend's worth. Paul suggested including this analysis code into the next offline monitoring run.
    • FDC - No news. Lubomir is still studying differences in alignment between Fall 2014 data and Spring 2015 data.
    • TOF - Brad has been able to run batch jobs over several runs and process the calibrations using Beni's scripts. The results still need to be reviewed.
    • SC - Eric has modified HDGeant to have paddle-by-paddle attenuation length and propogation time corrections, the values of which will be the same as used in the reconstruction. The results of these changes need to be studied in detail. Mahmoud is learning the current calibration codes and is getting ready to use them for future runs.
    • TAGH - Work still ongoing - timewalk corrections are next.
    • TAGM - Alex is working on the timewalk corrections. Elton pointed out that the scheme of obtaining timewalk corrections by plotting the TDC time versus the ADC peak won't work as well for the tagger detectors. Alex and Richard decided to calibrate against the coarse PS counters. Alex is working on refining the calibrations for the PSC to use as a benchmark.